Doug will make two $8000 investments. If one account yields 11%, what should the interest rate be on the other account to make a yearly total of $1720 in interest?
Answer by macston(5194)

I=Interest on second account
Subtract interest paid at 11% from total interest to get interest on second account only. Divide this by second account amount and multiply by 100% to get percent interest.
(($1720-($8000*.11))/$8000)(100%)=I
(($1720-$880)/$8000)(100%)=I
($840/$8000)(100%)=I
(0.105)(100%)=I
10.5%=I
ANSWER: The interest rate on the second account should be 10.5% to achieve a total return of $1720 on the investments.